if f is continuous for all x and, ∫ f(x)dx and ∫b/a f(x) dx are equal which one of the following statement is correct, 1. the function f is the constant e ,2. the function f is not differentiable, 3. the function f is the constant 0 , 4. the function f is any arbitrary constant c ,5. the function f is the identity function


1
Expert's answer
2021-05-28T10:18:22-0400

We have given the statement that,


"f" is continuous for all x


"\int f(x)dx = \dfrac{b}{a}\int f(x)dx"


Then, the function "f" is the constant